A boulder rolls off a 100m high cliff at a speed of 5m/s. How long will it take for the boulder to hit the ground? How far from the base of the cliff will the boulder land?

1
Expert's answer
2022-05-02T13:41:45-0400
"t=\sqrt{2H/g}=\sqrt{2*100/9.8}=4.5\:\rm s"

"L=v_0t=5*4.5=23\:\rm m"